# Beyond the Neural Fog: Interpretable Learning for AC Optimal Power Flow

### v1

@@ IV Comparison
In this section, we describe how we compare and evaluate the performance of the methods described in the previous section against the solution of the AC-OPF benchmark. The output of each approach is characterized by the active power generation levels ($\hat{\mathbf{p}}$) and the voltage magnitudes at the buses where these generators are located ($\hat{\mathbf{v}}$), while the true solution provided by the AC-OPF is denoted as $\mathbf{p}^{*}$ and $\mathbf{v}^{*}$. Since the DC-OPF model does not account for voltage magnitudes, they are assumed to be 1 p.u. for all buses. Similarly, the KT-DC approach, which is akin to the DC-OPF model (5), also omits voltage magnitudes. For this method, though, the voltage magnitudes at generating nodes are set to the values provided by the KNN regression that it uses. For each of the five methods, we compute the following three metrics:
@@
  • Infeasibility: This metric assesses the infeasibility of the obtained solutions. For each demand profile, active and reactive loads are fixed in demand nodes, the active output levels of all generators except the slack unit are fixed to those obtained by each approach, and the voltage magnitude levels are also fixed to the obtained solution. Subsequently, a standard AC power flow calculation is performed to determine an operating point that meets the power balance and power flow equations. We then assess the average normalized violations of the inequality constraints, following the methodology outlined in [22].
@@
    Note that model (6) is just a norm-1 projection into the AC-feasible space of the AC-OPF problem. Additionally, we provide the increase in operating cost of this AC-feasible point compared to the solution of the AC-OPF, which we denote as $\Delta C$ and compute as:
@@ V Numerical simulations
Table I presents a comprehensive overview of the performance metrics, as detailed in Section IV, averaged across 2 00020002\,0002 000 test instances for each network size. In this table, the metrics $MSE(\mathbf{p})$, $MSE(\mathbf{v})$, $\Delta\mathbf{p}$, and $\Delta\mathbf{v}$ are presented in per unit (p.u.), while the constraint violations and $\Delta C$ are expressed as percentages. To ease the interpretation of the results, the best performance that is achieved for each metric and network size is highlighted in bold.
@@
A comparison of the learning-based approaches shows that while NN generally surpasses KNN across the four systems, the differences are often minor. Particularly, for the 30- and 118-bus networks, KNN even yields lower values for $VIOL(\mathbf{p})$ and $VIOL(\mathbf{v})$. Therefore, although KNN may result in a reduction in performance due to its simplicity, it offers significant advantages, including lower data and computational requirements, as well as greater interpretability and explainability of the results.
Finally, comparing KT-DC and KT-AC reveals that the latter generally performs better, especially in approximating reactive power generation. While KT-AC is also a linear model, it accounts for both reactive flows and voltage magnitudes, resulting in more accurate outcomes. Despite this, KT-AC retains the advantages of a linear problem, offering convergence and optimality guarantees while being as efficiently solvable as the conventional DC-OPF problem. Additionally, the results show that KT-AC outperforms NN in most performance metrics across all network sizes. The only exceptions are the MSE of voltage magnitudes for the 30- and 118-bus networks, where NN performs slightly better. In all other cases, KT-AC delivers significantly superior performance, with improvements sometimes reaching one or two orders of magnitude. Thus, apart from being much simpler, more interpretable, and easier to train than neural networks, the proposed KT-AC provides solutions that more accurately approximate the true solution of the AC-OPF problem.
@@
Another limitation of neural networks is their high data dependency. To explore this, we conducted simulations by reducing the training data from 8 00080008\,0008 000 to 2 00020002\,0002 000 samples. The results for selected approaches are summarized in Table III. Comparing these results with those in Table I, we observe that the performance of KNN and KT-AC has remained relatively stable, whereas NN shows a significant degradation in performance. Specifically, for the reduced training set, KT-AC consistently outperforms NN across all analyzed metrics and network configurations. This highlights the superior data efficiency of the proposed learning-based method since KT-AC not only offers advantages in terms of speed, simplicity, and interpretability but also larger resilience to data scarcity, making it more practical for real-world applications.
@@ VI Conclusions
In this paper, we argue that while learning-based approaches can help improve the accuracy of power system solutions, it is essential to assign them a role compatible with a sector as critical and fundamental as the electrical industry. Only by recognizing and assuming the specificities of the operation of power systems, and the risk-averse mindset of their operators, we will be able to take the intermediate steps necessary for the seamless integration of machine learning and AI into existing power system operational routines. To support this point, we propose a hybrid strategy that effectively bridges the gap between model-based and learning-based methods. This strategy combines nearest neighbor regression with Taylor expansion, maintaining the simplicity of solving a linear optimization problem, similar to current model-based methods, while also improving accuracy by incorporating historical data. We introduce two versions of our proposed methodology. The first version, akin to the DC-OPF, focuses solely on the linear relationship between active power flows and voltage angles, omitting reactive power and voltage magnitudes. The second version, more aligned with the AC-OPF, incorporates both active and reactive power flows.

### v2

@@ IV Comparison
In this section, we describe how we compare and evaluate the performance of the methods described in the previous section against the solution of the AC-OPF benchmark. The output of each approach is characterized by the active power generation levels ($\hat{\mathbf{p}}$) and the voltage magnitudes at the buses where these generators are located ($\hat{\mathbf{v}}$), while the true solution provided by the AC-OPF is denoted as $\mathbf{p}^{*}$ and $\mathbf{v}^{*}$. Since the DC-OPF model does not account for voltage magnitudes, they are assumed to be 1 p.u. for all buses. Similarly, the KT-DC approach, which is akin to the DC-OPF model (5), also omits voltage magnitudes. For this method, though, the voltage magnitudes at generating nodes are set to the values provided by the KNN regression that it uses.
We initially evaluate the performance of the different methods using the mean squared error (MSE) of the active power generation levels and voltage magnitudes at the generation buses. However, since a lower MSE does not necessarily imply lower suboptimality, we also assess the methods using additional metrics that quantify the infeasibility and suboptimality of the obtained solutions. For each of the five methods, we compute the following three metrics:
@@
[…]
@@
    Note that model (6) is just a norm-1 projection into the AC-feasible space of the AC-OPF problem. In practice, if the slack variables are sufficiently small, these infeasibilities can often be accomodated by the system through minor violations of line capacity limits or slight deviations in system frequency. However, if the slack variables exceed acceptable thresholds, redispatch actions may be necessary to restore system feasibility and maintain operational reliability.
    Additionally, we provide the increase in operating cost of this AC-feasible point compared to the solution of the AC-OPF, as a measure of suboptimality, which we denote as $\Delta C$ and compute as:
@@ V Numerical simulations
Table I presents a comprehensive overview of the performance metrics, as detailed in Section IV, averaged across 2 00020002\,0002 000 test instances for each network size. In this table, the metrics $MSE(\mathbf{p})$, $MSE(\mathbf{v})$, $\Delta\mathbf{p}$, and $\Delta\mathbf{v}$ are presented in per unit (p.u.), while the constraint violations and $\Delta C$ are expressed as percentages. Besides, $\Delta\mathbf{p}$ and $\Delta\mathbf{v}$ are average values over the number of generators and the number of buses with generation, respectively. To ease the interpretation of the results, the best performance that is achieved for each metric and network size is highlighted in bold.
@@
A comparison of the learning-based approaches shows that while NN generally surpasses KNN across the four systems, the differences are often minor. Particularly, for the 30- and 118-bus networks, KNN even yields lower values for $VIOL(\mathbf{p})$ and $VIOL(\mathbf{v})$. Therefore, although KNN may result in a reduction in performance due to its simplicity, it offers significant advantages, including lower data and computational requirements, as well as greater interpretability and explainability of the results. Furthermore, the low MSE values for the KNN approach validate the use of this solution as the operating point for the first-order Taylor expansion.
Finally, comparing KT-DC and KT-AC reveals that the latter generally performs better, especially in approximating reactive power generation. While KT-AC also includes a linear formulation of the reactive power flow equations, it accounts for both reactive flows and voltage magnitudes, resulting in more accurate outcomes. Despite this, KT-AC retains the advantages of a convex problem, offering convergence and optimality guarantees while being as efficiently solvable as the conventional DC-OPF problem. Additionally, the results show that KT-AC outperforms NN in most performance metrics across all network sizes. The only exceptions are the MSE of voltage magnitudes for the 30- and 118-bus networks, where NN performs slightly better. In all other cases, KT-AC delivers significantly superior performance, with improvements sometimes reaching one or two orders of magnitude. Thus, apart from being much simpler, more interpretable, and easier to train than neural networks, the proposed KT-AC provides solutions that more accurately approximate the true solution of the AC-OPF problem.
@@
The data-driven approaches compared in this study may be influenced by the size of the training dataset. To investigate this, Table III presents the computational results obtained when the training sample size is reduced from 8 00080008\,0008 000 to 2 00020002\,0002 000 samples. Comparing these results with those in Table I, we observe that the performance of KNN and KT-AC has remained relatively stable, whereas NN shows a significant degradation in performance. Specifically, for the reduced training set, KT-AC consistently outperforms NN across all analyzed metrics and network configurations. This highlights the superior data efficiency of the proposed learning-based method since KT-AC not only offers advantages in terms of speed, simplicity, and interpretability but also larger resilience to data scarcity, making it more practical for real-world applications.
@@
To conclude this case study, we examine the potential of using solutions from learning-based approaches as starting points to reduce the number of iterations in the conventional AC-OPF model. Table IV reports the average number of iterations for 2 00020002\,0002 000 test samples of the 118-bus system, considering training datasets of either 8 00080008\,0008 000 or 2 00020002\,0002 000 samples for the learning approaches KNN, NN, and KT-AC. Additionally, we compare the iterations when the starting point is either the direct (and likely AC-infeasible) solution of each learning method or the corresponding feasible projection obtained through the post-processing model (6). With an average of 26.3 iterations for the AC-OPF model, the results reveal the following: for the larger training dataset, all learning methods reduce the average number of iterations. However, for the smaller training dataset, directly using NN solutions significantly increases iterations, highlighting the critical dependence of this approach on high-quality data. Besides, processing learning-based solutions to ensure feasibility consistently reduces iterations compared to using raw outputs. Importantly, these results show that the proposed KT-AC approach achieves the lowest average number of iterations among all methods.
TABLE IV: Effect of warm start on the average number of iterations of the AC-OPF model for the 118-bus system
8 00080008\,0008 000 training 2 00020002\,0002 000 training
INFES FEAS INFES FEAS
KNN 22.5 21.7 22.0 20.9
NN 22.1 21.3 517.3 22.4
KT-AC 20.8 20.4 20.8 20.3
@@ VI Conclusions
In this paper, we argue that while learning-based approaches can help improve the accuracy of power system solutions, it is essential to assign them a role compatible with a sector as critical and fundamental as the electrical industry. Only by recognizing and assuming the specificities of the operation of power systems, and the risk-averse mindset of their operators, we will be able to take the intermediate steps necessary for the seamless integration of machine learning and AI into existing power system operational routines. To support this point, we propose a hybrid strategy that effectively bridges the gap between model-based and learning-based methods. This strategy combines nearest neighbor regression with Taylor expansion, maintaining the simplicity of solving a convex optimization problem, similar to current model-based methods, while also improving accuracy by incorporating historical data. We introduce two versions of our proposed methodology. The first version, akin to the DC-OPF, focuses solely on the linear relationship between active power flows and voltage angles, omitting reactive power and voltage magnitudes. The second version, more aligned with the AC-OPF, incorporates both active and reactive power flows. The proposed method bridges the gap between traditional white-box approximations, such as DC-OPF, and modern black-box neural networks by providing a practical, interpretable, and accurate solution to the AC-OPF problem. Rather than relying entirely on black-box learning-based approaches to directly solve the AC-OPF, our method leverages historical data and learning techniques in a complementary role, preserving the interpretability and transparency of the resulting solution.
@@
Exploring the extension of the proposed approach to scenarios without historical optimal solution data, through the use of self-supervised methodologies, is identified as a promising avenue for future research. Another direction for future research is to extend the proposed algorithm to generate a distribution of dispatch decisions. This can be achieved by performing a Taylor expansion for each of the closest neighbors and solving multiple convex optimization problems, one for each neighbor.